OpenStudy (anonymous):

Find the common ratio of the infinite geometric series with the given sum and first term S=18 , a (sub) 1 =12

hartnn (hartnn):

use the formula \(\huge S_\infty = \dfrac{a_1}{r-1} \)

OpenStudy (anonymous):

i did but i cant get past the step to divide both sides by 10 .... i have examples in my book but its not varey informative

hartnn (hartnn):

i mean 1-r 18 = 12/(1-r) so,1-r = (12/18) can you proceed?

OpenStudy (anonymous):

yeah i got to the point where 1-r = 2/3